# DeepPolar: Inventing Nonlinear Large-Kernel Polar Codes via Deep Learning

### v1

@@ C.2 DeepPolar-parallel decoder
A major drawback of SC decoding and its variants like SCL decoders is the latency overhead due to sequential decoding. This drawback is inherently present in the DeepPolar-SC decoder. One method to break this latency barrier, whilst reducing the computational complexity of the decoder is to replace the $\ell$ SC-style NN decoders for a kernel at the lowest depth, by a single 3-layer FCNN, with hidden size 64, that decodes $\ell$ bits in one shot. We refer to this as DeepPolar-parallel. This resulted in an immediate reduction of 8x in parameter count compared to DeepPolar-SC while achieving better reliability, as highlighted in Fig. 14.
DeepPolarPolarDeepPolar-(parallel)SNR (dB)
@@ Appendix D Robustness to non-AWGN channels
Traditionally, the theoretical study of code design has been limited to AWGN channels. This is because of the difficulty in achieving a closed-form expression for the performance of the code in the presence of more complicated channel or noise conditions. But in practice, the channel experienced is far from AWGN, so testing any code on such varying channel conditions is crucial. We are specifically interested in the robustness of DeepPolar codes designs for AWGN channel, against non-AWGN channels .ie., we test and compare the performance without retaining.
DeepPolarBERPolar BERDeepPolarBLERPolar BLERSignal-to-noise ratio (SNR) [dB]
@@
(a) $n=256,k=37$, Rayleigh fast fading channel
DeepPolarBERPolar BERDeepPolarBLERPolar BLERSignal-to-noise ratio (SNR) [dB]
@@
(b) $n=256,k=37$, Bursty noise channel
Figure 15: Robustness to channel variations: DeepPolar (256,37,$\ell=16$) trained on AWGN channel maintains the gains over KO(8,2) as well as RM(8,2) and Polar(8,2) when tested on Rayliegh fast fading and bursty noise channels.
@@
where $h\sim\mathcal{N}(0,1)$ is the fading coefficient and $z_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the Gaussian noise. Note that the fading coefficient changes for every symbol in the codeword, making it a much worse channel than the AWGN channel. As seen from  Fig. 15a, the DeepPolar code demonstrates a gain of up to 0.3 dB compared to Polar(256,37), demonstrating the robustness of the learned code.
Next, we consider a bursty noise channel. Bursty noise, also known as popcorn noise, is a common type of noise encountered in semi-conductors. It can be described mathematically as
@@
where $z_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the Gaussian noise and $w_{i}\sim\mathcal{N}(0,\sigma^{2}_{b})$ with probability $\rho$ and $w_{i}=0$ with probability $1-\rho$ is the bursty noise. For our experiment, we choose $\rho=0.1$ and $\sigma_{b}=\sqrt{2}\sigma$. As seen from  Fig. 15b, the DeepPolar code demonstrates a gain of up to 0.25 dB compared to Polar(256,37) in the presence of bursty noise.
@@ Appendix E Experimental details
The complete source code is provided at: https://anonymous.4open.science/r/deeppolar-ED52
@@ E.2 Training SNR and number of steps
E.3 Large batch size and gradient accumulation
@@ E.4 Hyperparameter choices for DeepPolar (256,37)
The hyperparamters used for training DeepPolar (256,37,$\ell=16$) are listed below.
@@ E.5.1 Encoder
Figure 16: Each kernel $g$ at the encoder is parameterized by a fully connected network of 3 hidden layers and a hidden dimension of 64. The features ncorresponding to the Plotkin transform of size $\ell$ is provided via a skip connection

### v2

@@ C.2 DeepPolar-parallel decoder
Figure 15: DeepPolar ($n$=16,$k$$\ell$=4) : DeepPolar-parallel decoder retains the same decoding structure as DeepPolar-SC, but uses a one-shot parallel decoder at the leaves.
A major drawback of SC decoding and its variants like SCL decoders is the latency overhead due to sequential decoding. This drawback is inherently present in the DeepPolar-SC decoder. One method to break this latency barrier, whilst reducing the computational complexity of the decoder is to replace the $\ell$ SC-style NN decoders for a kernel at the lowest depth, by a single 3-layer FCNN, with hidden size 64, that decodes $\ell$ bits in one shot, as depicted in Fig. 15. We refer to this as DeepPolar-parallel. This resulted in an immediate reduction of 8x in parameter count compared to DeepPolar-SC while achieving the same reliability, as highlighted in Fig. 16.
DeepPolar-SCDeepPolar-parallelPolar-SCSNR (dB)
@@ Appendix D Robustness to non-AWGN channels
Traditionally, the canonical setting for the design and evaluation of codes has predominantly focused on AWGN channels due to the relative simplicity of closed-form performance analysis. However, practical communication channels often deviate significantly from the AWGN model. Ideally, both the code and decoder should be robust to various channels or adaptable to the encountered channel conditions. In this section, we demonstrate that DeepPolar codes, which are trained on AWGN channels, exhibit robustness when tested on non-AWGN channels without retraining. Further, in cases where the target channel model significantly diverges from AWGN noise, data-driven codes like DeepPolar can achieve substantial performance improvements through fine-tuning with a small number of steps on the target channel.
DeepPolarBERPolar BERDeepPolarBLERPolar BLERSNR (dB)
@@
(a)) ($n=256,k=37$), Rayleigh fast fading channel
DeepPolarBERPolar BERDeepPolarBLERPolar BLERSNR (dB)
@@
(b)) ($n=256,k=37$), Bursty noise channel
Figure 17: Robustness to channel variations: DeepPolar ($16256,37,\ell=$ 16) trained on AWGN channel maintains the gains over $\text{Polar}(256,37)$ when tested on Rayliegh fast fading and bursty noise channels.
SNR (dB)
Bit Error Rate
(a)) ($n=256,k=37$), Rayleigh fast fading channel
SNR (dB)
Bit Error Rate
(b)) ($n=256,k=37$), Bursty noise channel
Figure 18: Adaptivity: Performance of DeepPolar ($16256,37,\ell=$ 16) can be improved by adapting it to target channels. On channels such as Rayleigh fast fading, we do not see extra gains over the base decoder trained on AWGN. However, on large bursty noise we see a pronounced improvement
@@
where $h\sim\mathcal{N}(0,1)$ is the fading coefficient and $z_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the Gaussian noise. Note that the fading coefficient changes for every symbol in the codeword, making it a much worse channel than the AWGN channel. As seen from  Fig. 17a), the DeepPolar code demonstrates a gain of up to 0.3 dB compared to $\text{Polar}(256,37)$, demonstrating the robustness of the learned code.
Next, we consider a bursty noise channel. Bursty noise, also known as popcorn noise, can be described mathematically as
@@
where $z_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the Gaussian noise and $w_{i}\sim\mathcal{N}(0,\sigma^{2}_{b})$ with probability $\rho$ and $w_{i}=0$ with probability $1-\rho$ is the bursty noise. For our experiment, we choose $\rho=0.1$ and $\sigma_{b}=\sqrt{2}\sigma$. As seen from  Fig. 17b), the DeepPolar code demonstrates a gain of up to 0.25 dB compared to $\text{Polar}(256,37)$ in the presence of bursty noise.
Under larger degradation, for instance a bursty channel with $\rho=0.05$ and burst power $\sigma_{b}$= $10$, DeepPolar maintains its robustness. Still, the performance can be greatly enhanced by fine-tuning the decoder specifically for the target channel model, as demonstrated in  Fig. 18b). This adaptivity to unseen channel models is a key advantage of neural codes and decoders, such as DeepPolar, compared to classical codes.
@@ Appendix E Experimental details
The complete source code is provided at: https://www.github.com/hebbarashwin/deeppolar
@@ E.2 Training SNR and number of steps
High SNR performance. Obtaining gains over classical codes in the high SNR regime is challenging - this stems from several factors : Firstly, the error events become exceedingly sparse in the high SNR regime, leading to noisy gradient estimates and unstable training. Secondly, in the high SNR regime, the error rate is primarily governed by the code’s minimum distance. While DeepPolar exhibits remarkable distance properties (closely resembling the Gaussian codebook - Fig. 7), optimizing the minimum distance is challenging given the exponentially large space of codewords.
Nevertheless, high SNR performance can be boosted by finetuning the trained models at substantially high batch sizes. For instance, we finetune the model using a batch size of 200K (a 10x increase from the training phase), at -1dB using a learning rate of $10^{-5}$.
E.3 Large batch size
@@ E.4 Hyperparameter choices for DeepPolar (256,37)
The hyperparamters used for training DeepPolar (256,37,$\ell=16$) are listed in Table 1.
@@ E.5.1 Encoder
Figure 19: Each kernel $g$ at the encoder is parameterized by a fully connected network of 3 hidden layers and a hidden dimension of 64. The features corresponding to the Plotkin transform of size $\ell$ is provided via a skip connection
@@ F.1 Block Error Rate
SNR (dB)
Error Rate
Figure 22: Finetuning a trained DeepPolar model on a BLER surrogate objective trades off BER gains for BLER.
The figures of merit for any channel code are the Bit Error Rate (BER), and Block Error Rate (BLER). Since these metrics are non-differentiable, we need to find a differentiable surrogate. While the binary cross-entropy loss (used in the paper) acts as a stable surrogate loss function to optimize the BER, identifying such a surrogate loss for the BLER is an unsolved problem. (Often, optimizing for BER results in good BLER (eg - Fig. 21b). However, this is not guaranteed).
As preliminary work, we consider several potential loss functions to optimize the BLER, including $L_{1}=1-\Pi_{i=1}^{n}\sigma(u_{k}L_{k})$, where $L_{k}$ is the logit corresponding to the kth bit, and $L_{2}=\text{LogSumExp}(x_{1}^{n})$, where $x_{k}=\text{BCE}_{k}$.
However, both these losses lead to unstable training. For example, the product loss $L_{1}$ tends to show vanishing gradients. Nevertheless, fine-tuning a model initially trained on BCE loss to optimize the BER, on BLER loss $L_{1}$, yields improved BLER performance, demonstrated in Fig. 22 This strategy, however, introduces a tradeoff in BER gains - since the BER and BLER objectives are not fully aligned. With the right training objectives, neural codes provide flexibility in optimizing the figures of merit that we are interested in.
F.2 Model capacity ablation
Enc hidden size # Parameters in Enc Dec hidden size # Parameters in Dec BER at SNR -1.0 dB
8 5K 8 16K 9e-3
16 12K 16 44K 7.4e-4
16 12K 32 133K 2.1e-4
32 33K 32 133K 8.3e-5
64 100K 128 1.6M 5.9e-5
Table 2: Performance comparison as a function of the number of parameters in the encoder and decoder.
SNR (dB)
Bit Error Rate
SNR (dB)
Block Error Rate
Figure 23: Performance as a function of model size with pretraining and curriculum training: DeepPolar ($16256,37,\ell=$ 16)
DeepPolar-SC consists of a total of 100K+1.6M parameters; each kernel is an NN of depth 3, with hidden sizes of 64 and 128, respectively, at the encoder and decoder. We use the notation $(e=64,d=128)$ to represent this model. Generally, overparameterized models are easier to train. Since model size is an important practical consideration, we study the effect of model size on reliability. We consider the encoder and decoder networks with kernels with hidden sizes $e$ and $d$ respectively with depth $3$. We train the networks via kernel pretraining and curriculum learning, as outlined in Sec. 3.2. As shown in Table 2, Fig. 23, there is a noticeable degradation in performance when small models, i.e., $(e=8,d=8)$, $(e=16,e=16)$, are used to parameterize each kernel. However, $(e=16,e=32)$, and specifically $(e=32,h=32)$ with 160K parameters (10x reduction), outperforms SC in terms of BER and approaches the performance of $(e=64,d=128)$. These observations can be attributed to the dimension of the input, which is the kernel size 16.
F.3 Fully connected neural code - lacks generalization.
Figure 24: Codes learned via Fully connected NNs fail to generalize to unseen messages due to the absence of strong inductive biases.
While it might seem intuitive that non-linear codes would give better results compared to linear codes, the construction of good non-linear codes is highly non-trivial. To illustrate the difficulty, we train a dense, fully connected encoder-decoder pair using the same training methodology as DeepPolar (Sec. E). As shown in Fig. 24, despite this network having ten times more parameters than DeepPolar, its performance is substantially inferior to the successive cancellation (SC) decoding of polar codes. This is an effect of the model’s poor generalization in the absence of good inductive biases, a phenomenon well-documented in existing literature, for example, see (Jiang et al., 2020b). Thus, neural architectures that add redundancy in a structured way are essential for learning a good code.